# The impact of $\gamma N$ and $\gamma^{*}N$ interactions on our understanding of nucleon excitations

### v1

@@ 5.2.2   The $\gamma^{*}p\to N(1440)\nicefrac{{1}}{{2}}^{+}$ transition and the Roper resonance
Figure 5.9: a: Trajectories of the evolution of the $J^{P}=\nicefrac{{1}}{{2}}^{+}$ resonance. Three poles evolve from one bare $N^{*}$ state at 1763 MeV [437]. b: Lattice results for the nucleon and Roper masses (adapted from Ref. [877]). Most data points for the Roper resonance lie above the $N\pi$ threshold. c and d: Lattice results for the $\gamma^{\ast}p\to N(1440)$ proton to Roper transition form factors $F_{1}$ and $F_{2}$ for pion masses 390 MeV (red squares), 450 MeV (orange triangles) and 875 MeV (green circles) compared to data (open circles) [933].
@@
Although the zero crossing in the transition form factor $F_{2}$ suggests an interpretation of the Roper resonance as the first radial excitation of the nucleon, the fact that it strongly couples to meson-baryon channels at physical pion masses also allows for different interpretations [883, 934]. Combining lattice QCD and Hamiltonian EFT, the authors of Ref. [935] argue that the 2s radial excitation of the nucleon should have a mass of 2 GeV, while the Roper resonance is generated by rescattering effects in the coupled channels of $\pi N$, $\sigma N$, and $\pi\Delta(1232)$. In a similar spirit, the coupled-channel analysis of Ref. [437] assumes one “bare" pole at 1763 MeV, which couples to $\pi N$, $\eta N$, $\pi\Delta$, $\rho N$ and $\sigma N$ and thereby generates three further poles (see Fig. 5.9 a). One is interpreted as the Roper resonance and another one as the $N(1710)\nicefrac{{1}}{{2}}^{+}$, suggesting that both resonances originate from the same excited nucleon state. This possibility is discussed in more detail in Sec. 5.2.8.
@@
Here we would like to emphasize again that, from the point of view of quantum field theory, the options “three-quark state” or “dynamically generated resonance” are not exclusive. As we argued in Sec. 4.2, if a baryon pole appears in the $qqq$ correlation function, it must also appear in the $qqqq\bar{q}$, $qqqq\bar{q}q\bar{q}$ and any other correlation function that can produce baryon poles. Unless some quantum number prevents it, there is no dynamically generated resonance without an accompanying quark core. This implies that there is also no contradiction between the Roper resonance being the radial excitation of the nucleon while at the same time having a large admixture of five- or even seven-quark components. A similar example in the meson sector is the famous $\chi_{c1}(3872)$, which couples to $c\bar{c}$ and $c\bar{c}q\bar{q}$ and is therefore simultaneously a $c\bar{c}$ excitation and a four-quark state [839, 936, 606].
@@
Similar conclusions about the nature of the Roper resonance have been reached in different approaches analyzing the $Q^{2}$ dependence of electroproduction data [937, 938, 939, 18]. As stated in the conclusions of [546]: "The fifty years of experience with the Roper resonance have delivered lessons that cannot be emphasized too strongly. Namely, in attempting to predict and explain the QCD spectrum, one must fully consider the impact of meson-baryon final state interactions and the coupling between channels and states that they generate, and look beyond merely locating the poles in the S-matrix, which themselves reveal little structural information, to also consider the $Q^{2}$ dependencies of the residues, which serve as a penetrating scale-dependent probe of resonance composition."
Figure 5.10: Helicity transition amplitudes $A_{1/2}$, $A_{3/2}$ and $S_{1/2}$ for the transition $\gamma^{*}p\to\Delta(1600)\nicefrac{{3}}{{2}}^{+}$. The data are from [289], the curves are from [784].
@@ 5.2.3   The $\gamma^{*}p\to\Delta(1600)\nicefrac{{3}}{{2}}^{+}$ transition
The recent results from the CLAS collaboration in $\gamma^{*}p\to p\pi^{+}\pi^{-}$ electro-production [289] have changed the situation significantly and enabled the first extraction of the $Q^{2}$ dependence of the transition amplitudes in the range $4.52<Q^{2}<$ 4.5 GeV2 as shown in Fig. 5.10. The theory curves are based on the DSE approach using a dynamical quark-diquark approximation to the relativistic three-body bound state problem. It provides an excellent fit to all three helicity amplitudes and establishes the resonance as a solid three-quark radial excitation. The lack of electroproduction data at low $Q^{2}$ prohibits an evaluation of possible meson-baryon contributions in the lower $Q^{2}$ range. It is worth noting that the theory predictions existed before the data became available.
An interesting question regarding the $\Delta(1600)\nicefrac{{3}}{{2}}^{+}$ as a radial excitation of the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$ is whether it behaves similarly to the $N(1440)\nicefrac{{1}}{{2}}^{+}$ as a radial excitation of the nucleon ground state, i.e., whether it shows a similar zero-crossing of its helicity amplitudes. This has been studied within the LF RQM [299]. In this work the configuration mixings that follow from QCD-inspired inter-quark forces on the results for the electroexcitation of the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$, $N(1440)\nicefrac{{1}}{{2}}^{+}$, and $\Delta(1600)\nicefrac{{3}}{{2}}^{+}$ obtained earlier in the light-front relativistic quark model without configuration mixing [314]. In the case of the $N(1440)\nicefrac{{1}}{{2}}^{+}$, the configuration mixing does improve the overall agreement with the data while not significantly changing the zero crossing of the transverse transition amplitude $A_{1/2}$. For the $\gamma^{\ast}N\to\Delta(1600)\nicefrac{{3}}{{2}}^{+}$ transition, configuration mixing changes very strongly the results obtained earlier for the $N$ and $\Delta(1600)\nicefrac{{3}}{{2}}^{+}$ taken as pure states in the multiplets $(\textbf{56},0^{+})$ and $(\textbf{56}^{\prime},0^{+})$. While in the case without mixing both amplitudes $A_{1/2}$ and $A_{3/2}$ are predicted to change sign at $Q^{2}\approx 0.25$ GeV2, both remain negative when mixing is taken into account, as also measured at the real photon point, with a smooth falloff in magnitude with increasing $Q^{2}$. The predicted behavior is consistent with the experimental data obtained at $Q^{2}>2~{}$GeV2 and shown in Fig. 5.10.
@@ 5.2.4   The $\gamma^{*}p\to N(1520)\nicefrac{{3}}{{2}}^{-}$ transition and switch of its helicity structure
[…]
@@
Figure 5.11: The $\gamma^{\ast}p\to N(1520)$ resonance helicity transition amplitudes $A_{1/2}$, $A_{3/2}$, $S_{1/2}$ and helicity transition asymmetry $A_{hel}$. Data and LF RQM results from [320].
A similar behavior is expected for the $N(1680)\nicefrac{{5}}{{2}}^{+}$ with a somewhat slower change in the helicity structure with $Q^{2}$ as seen in the comparison in Eq. (5.2). A somewhat analogous analysis has been conducted by Ramalho and Pena [940] within a covariant spectator quark-diquark model that assumes additional pion cloud contributions. A significant difference to the approach adopted here is that the $A_{3/2}$ helicity amplitude does not have a quark core contribution but coincides with the pion cloud contribution. Therefore, their results cannot be compared directly with the approach taken here.
@@ 5.2.5    The $\gamma^{*}p\to N(1535)\nicefrac{{1}}{{2}}^{+}$ transition
The $N(1535)\nicefrac{{1}}{{2}}^{+}$ is the parity partner of the ground-state nucleon, with the same spin $J=1/2$ but with opposite parity. Its quark content requires an orbital $L=1$ excitation in the transition from the proton. In the $SU(6)\otimes O(3)$ symmetry scheme, the state is a member of the $(\textbf{70},1^{-})$ supermultiplet and couples nearly equally to the $N\pi$ and $N\eta$ final states. It has therefore been probed using both decay channels in $ep\to ep\eta$ and $ep\to eN\pi$, which are shown in Fig. 5.13. For isospin $I=1/2$ nucleon states, the coupling to the charged $\pi^{+}n$ channel is preferred over $\pi^{0}p$ owing to the Clebsch-Gordon coefficients. The $A_{1/2}$ helicity amplitude for the $\gamma^{*}p\to N(1535)\nicefrac{{1}}{{2}}^{-}$ resonance excitation shown in Fig. 5.13 is known over the widest range in $Q^{2}$ of all nucleon states for which resonance transition form factors have been measured as part of the broad experimental program at JLab. We therefore may draw more definitive conclusions compared to some other states. For this resonance, as well as for $N(1440)\nicefrac{{1}}{{2}}^{+}$, advanced relativistic quark model calculations [314], DSE calculations [781] and Light Cone sum rule results [914] are available, employing QCD-based modeling of the excitation of the quark core for the first time.
@@
The transverse transition amplitude $A_{1/2}$ of $N(1535)\nicefrac{{1}}{{2}}^{-}$ is a prime example of the power of meson electroproduction to unravel the internal structure of the resonance transition. The nature of the $N(1535)\nicefrac{{1}}{{2}}^{-}$ resonance has been discussed as a dynamically generated resonance, e.g. [941, 635]. The electroproduction data discussed here reveal structural aspects of the state and its nature that support a different interpretation. The transition form factor $A_{1/2}$ of the state, shown in Fig. 5.13, is quantitatively reproduced over a large range in $Q^{2}$ by two alternative approaches, the LF RQM and the LCSR. Both calculations are based on the assumption of the presence of a three-quark core. Note that there is a significant deviation from the quark calculations essentially only at $Q^{2}<1\dots 2$ GeV2 for both three-quark core related calculations. The difference to the data is highlighted for the LF RQM calculation as the shaded area in Fig. 5.13, which may be assigned to the presence of non-quark contributions. We also note that the results of a multipole analysis shown in Fig. 5.13 qualitatively confirms the conclusions drawn from the discussion above. Attempts to compute the transition form factors within a strictly dynamical model, shown as the DGR labeled curve, have not succeeded in explaining the data in any part of the $Q^{2}$ range covered [917]. However, in this case only the modulos of $A_{1/2}$ is shown in the DGR curve, and the phase of the amplitude was adjusted to obtain a value > 0. We may therefore change the sign of that contribution, which then would make the DGR line closely resemble the shaded MB contributions, both in sign and in absolute value at small $Q^{2}$. We may then conclude that the DGR result is dominated by the meson-baryon cloud contribution, which is not included in either of the three-quark core related calculations. Small meson-baryon contributions are also consistent with [942] who found negligible compositeness for the $N(1535)\nicefrac{{1}}{{2}}^{-}$ state. More generally, Ref. [943] found that near-threshold resonances have small composite fractions, which applies to the $N(1535)\nicefrac{{1}}{{2}}^{-}$ with a pole position of 1510–1520 MeV, just above the $\eta$-production threshold.
@@
A caveat arises when translating the helicity amplitudes to constraint-free transition form factors $F_{1}$ and $F_{2}$. While many models reproduce the qualitative behavior of $F_{1}$, they struggle to describe $F_{2}$ which drops unusually fast [18]. In the EBAC analysis [436] as well as the covariant spectator model [700] the ‘bare’ contribution to $F_{2}$ even has an opposite sign, which points towards substantial meson cloud contributions. On the other hand, models that include such contributions describe the data well [944, 945].
@@
Figure 5.12: (Color online) $A_{1/2}$ transition amplitude. The data include both $N\pi$ and $p\eta$ in the final state. The dashed line is the result of assuming $N(1535)$ is a dynamically generated resonance (DGR) and applying a form factor that is external to the model. [917]
@@
In previous discussions we have concluded that modern quark-based models, tuned to reproduce the high-$Q^{2}$ behavior of the electroproduction transition amplitudes $A_{1/2}$, $A_{3/2}$, fail to reproduce quantitatively the behavior in $Q^{2}$ ranges below 1 …2 GeV2. In this range meson-baryon degrees of freedom provide significant strength to the resonance excitation depending on the specific character of the excited state. For the Roper resonance, the effective field theory approach of Ref. [922] reproduces the transition amplitudes at $Q^{2}<0.6$ GeV2, whereas quark-based approaches such as the LF RQM and the DSE and LCSR calculations do not reproduce the transition amplitudes quantitatively. Nevertheless, the LF RQM and DSE approaches predict features like the change of sign of the $A_{1/2}$ amplitude as seen in Fig. 5.8, but at somewhat lower $Q^{2}$.
@@
In the case of the $N(1675)\nicefrac{{5}}{{2}}^{-}$, we can actually measure the meson-baryon contributions directly without having to separate them from the quark core contribution. This is due to the suppression of the quark contributions in the excitation of this resonance when probed on proton targets. Fig. 5.14 shows our current knowledge of the transverse helicity amplitudes $A_{1/2}(Q^{2})$ and $A_{3/2}(Q^{2})$ for a proton target, compared to LF RQM [298] and hypercentral constituent-quark model [528] calculations. The specific quark transition for a $J^{P}=\nicefrac{{5}}{{2}}^{-}$ state, belonging to the $[SU(6)\otimes O(3)]=(\mathbf{70},1^{-})$ supermultiplet configuration, in the non-relativistic approximation prohibits the transition from the proton in a single quark transition.
@@
Figure 5.14: (Color online) Left and center: Transverse helicity amplitudes $A_{1/2}$ and $A_{3/2}$ for $\gamma^{*}p\to N(1675)\nicefrac{{5}}{{2}}^{-}$ (in $10^{-3}$ GeV-1/2). Right: The amplitude $A_{1/2}$ on neutron target. The solid curves are from the LF RQM and the hypercentral constituent quark model (hCQM) [528]. The dashed-dotted curves are absolute values of the predicted meson-baryon contributions from the dynamical coupled-channel approach of Ref. [946]. The solid circles are the amplitudes extracted from CLAS pion electroproduction data [320]. The solid square at $Q^{2}=0$ is the RPP estimate [947]. Note that, as an effect of the selection rule, the projected quark model amplitude $A_{1/2}$ of the neutron at the photon point is one order of magnitude larger than the one on the proton.
@@
This suppression, known as the Moorhouse selection rule [948], is valid for the transverse transition amplitudes $A_{1/2}$ and $A_{3/2}$ at all $Q^{2}$. It applies to the transition from a proton target but not from the neutron, which is consistent with the data at the $Q^{2}=0$ point, where data from both proton and neutron targets exist. Modern relativistic quark models that go beyond single-quark transitions quantitatively confirm this suppression, resulting in very small amplitudes from protons but large ones from neutrons. Furthermore, a direct computation of the hadronic contribution to the transition from protons in a coupled-channel approach confirms this (Fig. 5.14). The measured helicity amplitudes off the protons are almost exclusively due to MB contributions, as the dynamical coupled-channel (DCC) calculation indicates (dashed line). The close correlation of the DCC calculation and the measured data for the case when quark contributions are nearly absent supports the phenomenological description in terms of a three-quark core that dominates at high $Q^{2}$ and MB contributions that can make important contributions at lower $Q^{2}$. From the direct measurement of the MB contributions on the proton, and from measurements on the neutron at the photon point subtracted of the LF RQM predictions, one finds that for this resonance the MB contributions are of isovector nature, as shown in Fig. 5.16.

### v4

@@ 5.2.2  The $\gamma^{*}p\to N(1440)\nicefrac{{1}}{{2}}^{+}$ transition and the Roper resonance
Figure 5.9: a: Trajectories of the evolution of the $J^{P}=\nicefrac{{1}}{{2}}^{+}$ resonance. Three poles evolve from one bare $N^{*}$ state at 1763 MeV Suzuki:2009nj . b: Lattice results for the nucleon and Roper masses (adapted from Ref. Liu:2016rwa ). Most data points for the Roper resonance lie above the $N\pi$ threshold. c and d: Lattice results for the $\gamma^{\ast}p\to N(1440)$ proton to Roper transition form factors $F_{1}$ and $F_{2}$ for pion masses 390 MeV (red squares), 450 MeV (orange triangles) and 875 MeV (green circles) compared to data (open circles) Lin:2011da .
@@
Although the zero crossing in the transition form factor $F_{2}$ suggests an interpretation of the Roper resonance as the first radial excitation of the nucleon, the fact that it strongly couples to meson-baryon channels at physical pion masses also allows for different interpretations Wu:2017qve , Virgili:2019shg . Combining lattice QCD and Hamiltonian EFT, the authors of Ref. Leinweber:2024psf argue that the 2s radial excitation of the nucleon should have a mass of 2 GeV, while the Roper resonance is generated by rescattering effects in the coupled channels of $\pi N$, $\sigma N$, and $\pi\Delta(1232)$. In a similar spirit, the coupled-channel analysis of Ref. Suzuki:2009nj assumes one “bare" pole at 1763 MeV, which couples to $\pi N$, $\eta N$, $\pi\Delta$, $\rho N$ and $\sigma N$ and thereby generates three further poles (see Fig. 5.9 a). One is interpreted as the Roper resonance and another one as the $N(1710)\nicefrac{{1}}{{2}}^{+}$, suggesting that both resonances originate from the same excited nucleon state. This possibility is discussed in more detail in Sec. 5.2.8.
@@
Here we would like to emphasize again that, from the point of view of quantum field theory, the options “three-quark state” or “dynamically generated resonance” are not exclusive. As we argued in Sec. 4.2, if a baryon pole appears in the qqqqqq correlation function, it must also appear in the $\bar{q}, qqqqq¯qq¯q$$\bar{q}q\bar{q}$ and any other correlation function that can produce baryon poles. Unless some quantum number prevents it, there is no dynamically generated resonance without an accompanying quark core. This implies that there is also no contradiction between the Roper resonance being the radial excitation of the nucleon while at the same time having a large admixture of five- or even seven-quark components. A similar example in the meson sector is the famous $\chi_{c1}(3872)$, which couples to $c\bar{c}$ and $c\bar{c}q\bar{q}$ and is therefore simultaneously a $c\bar{c}$ excitation and a four-quark state Prelovsek:2013cra , Padmanath:2015era , Guo:2017jvc .
@@
Similar conclusions about the nature of the Roper resonance have been reached in different approaches analyzing the $Q^{2}$ dependence of electroproduction data Karapetyan:2023kzs , Kaewsnod:2021nfw , Ramalho:2010js , Ramalho:2023hqd . As stated in the conclusions of Burkert:2017djo : "The fifty years of experience with the Roper resonance have delivered lessons that cannot be emphasized too strongly. Namely, in attempting to predict and explain the QCD spectrum, one must fully consider the impact of meson-baryon final state interactions and the coupling between channels and states that they generate, and look beyond merely locating the poles in the S-matrix, which themselves reveal little structural information, to also consider the $Q^{2}$ dependencies of the residues, which serve as a penetrating scale-dependent probe of resonance composition."
Figure 5.10: Helicity transition amplitudes $A_{1/2}$, $A_{3/2}$ and $S_{1/2}$ for the transition $\gamma^{*}p\to\Delta(1600)\nicefrac{{3}}{{2}}^{+}$. The data are from Mokeev:2023zhq , the curves are from Lu:2019bjs .
@@ 5.2.3  The $\gamma^{*}p\to\Delta(1600)\nicefrac{{3}}{{2}}^{+}$ transition
The recent results from the CLAS collaboration in $\gamma^{*}p\to p\pi^{+}\pi^{-}$ electro-production Mokeev:2023zhq have changed the situation significantly and enabled the first extraction of the $Q^{2}$ dependence of the transition amplitudes in the range $2<Q^{2}<4.5$ GeV2 as shown in Fig. 5.10. The theory curves are based on the DSE approach using a dynamical quark-diquark approximation to the relativistic three-body bound state problem. It provides an excellent fit to all three helicity amplitudes and establishes the resonance as a solid three-quark radial excitation. The lack of electroproduction data at low $Q^{2}$ prohibits an evaluation of possible meson-baryon contributions in the lower $Q^{2}$ range. It is worth noting that the theory predictions existed before the data became available.
An interesting question regarding the $\Delta(1600)\nicefrac{{3}}{{2}}^{+}$ as a radial excitation of the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$ is whether it behaves similarly to the $N(1440)\nicefrac{{1}}{{2}}^{+}$ as a radial excitation of the nucleon ground state, i.e., whether it shows a similar zero-crossing of its helicity amplitudes. This has been studied within the LF RQM Aznauryan:2016wwm . In this work the configuration mixings that follow from QCD-inspired inter-quark forces on the results for the electroexcitation of the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$, $N(1440)\nicefrac{{1}}{{2}}^{+}$, and $\Delta(1600)\nicefrac{{3}}{{2}}^{+}$ obtained earlier in the light-front relativistic quark model without configuration mixing Aznauryan:2015zta . In the case of the $N(1440)\nicefrac{{1}}{{2}}^{+}$, the configuration mixing does improve the overall agreement with the data while not significantly changing the zero crossing of the transverse transition amplitude $A_{1/2}$. For the $\gamma^{\ast}N\to\Delta(1600)\nicefrac{{3}}{{2}}^{+}$ transition, configuration mixing changes very strongly the results obtained earlier for the $N$ and $\Delta(1600)\nicefrac{{3}}{{2}}^{+}$ taken as pure states in the multiplets $(\textbf{56},0^{+})$ and $(\textbf{56}^{\prime},0^{+})$. While in the case without mixing both amplitudes $A_{1/2}$ and $A_{3/2}$ are predicted to change sign at $Q^{2}\approx 0.25$ GeV2, both remain negative when mixing is taken into account, as also measured at the real photon point, with a smooth falloff in magnitude with increasing $Q^{2}$. The predicted behavior is consistent with the experimental data obtained at $Q^{2}>2$~GeV2 and shown in Fig. 5.10.
@@ 5.2.4  The $\gamma^{*}p\to N(1520)\nicefrac{{3}}{{2}}^{-}$ transition and switch of its helicity structure
The $N(1520)\nicefrac{{3}}{{2}}^{-}$ state is the lowest excited nucleon resonance with $J^{P}=\nicefrac{{3}}{{2}}^{-}$. Its helicity structure is defined by the $Q^{2}$ dependence of the two transverse transition amplitudes $A_{1/2}$ and $A_{3/2}$, which are shown in Fig. 5.11. A particularly interesting feature of this state is that at the real-photon point $A_{3/2}$ is strongly dominant, while $A_{1/2}$ is very small. However, at increasing $Q^{2}$, $A_{1/2}$ becomes dominant, while $A_{3/2}$ drops rapidly. Meson-baryon contributions are small at $Q^{2}>1.0\dots 1.5$ GeV2 making the quark core excitation the dominant contribution. At $Q^{2}>2$~ GeV2, $A_{1/2}$ fully dominates the process. The helicity asymmetry $A_{hel}=(A_{1/2}^{2}-A_{3/2}^{2})/(A_{1/2}^{2}+A_{3/2}^{2})$ shown in the right panel of Fig. 5.11 illustrates this rapid change in the helicity structure. Expressed in terms of constraint-free form factors, the effect can be explained by the fact that two form factors have the same asymptotic falloff ($\propto 1/Q^{6}$) and cancel in the helicity amplitude $A_{3/2}$ but add up in $A_{1/2}$ Eichmann:2018ytt . Within the constituent-quark model picture, this behavior can be understood from the quark wave function, see Eq. (5.2) below, which on the one hand allows for $A_{1/2}=0$ at ${\mbox{$Q$}}^{2}=\alpha^{2}$, where $\alpha$ is the harmonic oscillator constant, and on the other hand causes a rapid rise of the ratio of $A_{1/2}/A_{3/2}$ with increasing $Q^{2}$.
@@
Figure 5.11: The $\gamma^{\ast}p\to N(1520)$ resonance helicity transition amplitudes $A_{1/2}$, $A_{3/2}$, $S_{1/2}$ and helicity transition asymmetry $A_{hel}$. Data and LF RQM results from CLAS:2009ces .
A similar behavior is expected for the $N(1680)\nicefrac{{5}}{{2}}^{+}$ with a somewhat slower change in the helicity structure with $Q^{2}$ as seen in the comparison in Eq. (5.2). A somewhat analogous analysis has been conducted by Ramalho and Pena Ramalho:2013mxa within a covariant spectator quark-diquark model that assumes additional pion cloud contributions. A significant difference to the approach adopted here is that the $A_{3/2}$ helicity amplitude does not have a quark core contribution but coincides with the pion cloud contribution. Therefore, their results cannot be compared directly with the approach taken here.
@@ 5.2.5   The $\gamma^{*}p\to N(1535)\nicefrac{{1}}{{2}}^{+}$ transition
The $N(1535)\nicefrac{{1}}{{2}}^{+}$ is the parity partner of the ground-state nucleon, with the same spin $J=1/2$ but with opposite parity. Its quark content requires an orbital $L=1$ excitation in the transition from the proton. In the $SU(6)\otimes O(3)$ symmetry scheme, the state is a member of the $(\textbf{70},1^{-})$ supermultiplet and couples nearly equally to the $N\pi$ and $N\eta$ final states. It has therefore been probed using both decay channels in $ep\to ep\eta$ and $ep\to eN\pi$, which are shown in Fig. 5.13. For isospin $I=1/2$ nucleon states, the coupling to the charged $\pi^{+}n$ channel is preferred over $\pi^{0}p$ owing to the Clebsch-Gordon coefficients. The $A_{1/2}$ helicity amplitude for the $\gamma^{*}p\to N(1535)\nicefrac{{1}}{{2}}^{-}$ resonance excitation shown in Fig. 5.13 is known over the widest range in $Q^{2}$ of all nucleon states for which resonance transition form factors have been measured as part of the broad experimental program at JLab. We therefore may draw more definitive conclusions compared to some other states. For this resonance, as well as for $N(1440)\nicefrac{{1}}{{2}}^{+}$, advanced relativistic quark model calculations Aznauryan:2015zta , DSE calculations Segovia:2015hra and Light Cone sum rule results Anikin:2015ita are available, employing QCD-based modeling of the excitation of the quark core for the first time.
@@
The transverse transition amplitude $A_{1/2}$ of $N(1535)\nicefrac{{1}}{{2}}^{-}$ is a prime example of the power of meson electroproduction to unravel the internal structure of the resonance transition. The nature of the $N(1535)\nicefrac{{1}}{{2}}^{-}$ resonance has been discussed as a dynamically generated resonance, e.g. Kaiser:1995cy , Bruns:2010sv . The electroproduction data discussed here reveal structural aspects of the state and its nature that support a different interpretation. The transition form factor $A_{1/2}$ of the state, shown in Fig. 5.13, is quantitatively reproduced over a large range in $Q^{2}$ by two alternative approaches, the LF RQM and the LCSR. Both calculations are based on the assumption of the presence of a three-quark core. Note that there is a significant deviation from the quark calculations essentially only at $Q^{2}<1\dots 2$ GeV2 for both three-quark core related calculations. The difference to the data is highlighted for the LF RQM calculation as the shaded area in Fig. 5.13, which may be assigned to the presence of non-quark contributions. We also note that the results of a multipole analysis shown in Fig. 5.13 qualitatively confirms the conclusions drawn from the discussion above. Attempts to compute the transition form factors within a strictly dynamical model, shown as the DGR labeled curve, have not succeeded in explaining the data in any part of the $Q^{2}$ range covered Jido:2007sm . However, in this case only the modulos of $A_{1/2}$ is shown in the DGR curve, and the phase of the amplitude was adjusted to obtain a value > 0. We may therefore change the sign of that contribution, which then would make the DGR line closely resemble the shaded MB contributions, both in sign and in absolute value at small $Q^{2}$. We may then conclude that the DGR result is dominated by the meson-baryon cloud contribution, which is not included in either of the three-quark core related calculations. Small meson-baryon contributions are also consistent with Sekihara:2015gvw who found negligible compositeness for the $N(1535)\nicefrac{{1}}{{2}}^{-}$ state. More generally, Ref. Kinugawa:2024kwb found that near-threshold resonances have small composite fractions, which applies to the $N(1535)\nicefrac{{1}}{{2}}^{-}$ with a pole position of 1510–1520 MeV, just above the $\eta$-production threshold.
@@
A caveat arises when translating the helicity amplitudes to constraint-free transition form factors $F_{1}$ and $F_{2}$. While many models reproduce the qualitative behavior of $F_{1}$, they struggle to describe $F_{2}$ which drops unusually fast Ramalho:2023hqd . In the EBAC analysis Julia-Diaz:2009dnz as well as the covariant spectator model Ramalho:2020nwk the ‘bare’ contribution to $F_{2}$ even has an opposite sign, which points towards substantial meson cloud contributions. On the other hand, models that include such contributions describe the data well Golli:2011jk , Gutsche:2019yoo .
@@
Figure 5.12: (Color online) $A_{1/2}$ transition amplitude. The data include both $N\pi$ and $p\eta$ in the final state. The dashed line is the result of assuming $N(1535)$ is a dynamically generated resonance (DGR) and applying a form factor that is external to the model. Jido:2007sm
@@
In previous discussions we have concluded that modern quark-based models, tuned to reproduce the high-$Q^{2}$ behavior of the electroproduction transition amplitudes $A_{1/2}$, $A_{3/2}$, fail to reproduce quantitatively the behavior in $Q^{2}$ ranges below 1 …2 GeV2. In this range meson-baryon degrees of freedom provide significant strength to the resonance excitation depending on the specific character of the excited state. For the Roper resonance, the effective field theory approach of Ref. Bauer:2014cqa reproduces the transition amplitudes at $Q^{2}<0.6$ GeV2, whereas quark-based approaches such as the LF RQM and the DSE and LCSR calculations do not reproduce the transition amplitudes quantitatively. Nevertheless, the LF RQM and DSE approaches predict features like the change of sign of the $A_{1/2}$ amplitude as seen in Fig. 5.8, but at somewhat lower $Q^{2}$.
@@
In the case of the $N(1675)\nicefrac{{5}}{{2}}^{-}$, we can actually measure the meson-baryon contributions directly without having to separate them from the quark core contribution. This is due to the suppression of the quark contributions in the excitation of this resonance when probed on proton targets. Fig. 5.14 shows our current knowledge of the transverse helicity amplitudes $A_{1/2}(Q^{2})$ and $A_{3/2}(Q^{2})$ for a proton target, compared to LF RQM Aznauryan:2017nkz and hypercentral constituent-quark model Santopinto:2012nq calculations. The specific quark transition for a $J^{P}=\nicefrac{{5}}{{2}}^{-}$ state, belonging to the $SU(6)\otimes O(3)]=(\mathbf{70},1^{-})$ supermultiplet configuration, in the non-relativistic approximation prohibits the transition from the proton in a single quark transition.
@@
Figure 5.14: (Color online) Left and center: Transverse helicity amplitudes $A_{1/2}$ and $A_{3/2}$ for $\gamma^{*}p\to N(1675)\nicefrac{{5}}{{2}}^{-}$ (in $10^{-3}$ GeV-1/2). Right: The amplitude $A_{1/2}$ on neutron target. The solid curves are from the LF RQM and the hypercentral constituent quark model (hCQM) Santopinto:2012nq . The dashed-dotted curves are absolute values of the predicted meson-baryon contributions from the dynamical coupled-channel approach of Ref. Julia-Diaz:2007mae . The solid circles are the amplitudes extracted from CLAS pion electroproduction data CLAS:2009ces . The solid square at $Q^{2}=0$ is the RPP estimate ParticleDataGroup:2016lqr . Note that, as an effect of the selection rule, the projected quark model amplitude $A_{1/2}$ of the neutron at the photon point is one order of magnitude larger than the one on the proton. The non-quark contributions for both amplitudes $A_{1/2}(0)$ and $A_{3/2}(0)$ are $15\pm 7\times 10^{-3}$GeV-1/2 for the proton, and $15\pm 7\times 10^{-3}$GeV-1/2 for the neutron.
@@
This suppression, known as the Moorhouse selection rule Moorhouse:1966jn , is valid for the transverse transition amplitudes $A_{1/2}$ and $A_{3/2}$ at all $Q^{2}$. It applies to the transition from a proton target but not from the neutron, which is consistent with the data at the $Q^{2}=0$ point, where data from both proton and neutron targets exist. Modern relativistic quark models that go beyond single-quark transitions quantitatively confirm this suppression, resulting in very small amplitudes from protons but large ones from neutrons. Furthermore, a direct computation of the hadronic contribution to the transition from protons in a coupled-channel approach confirms this (Fig. 5.14). The measured helicity amplitudes off the protons are almost exclusively due to MB contributions, as the dynamical coupled-channel (DCC) calculation indicates (dashed line). The close correlation of the DCC calculation and the measured data for the case when quark contributions are nearly absent supports the phenomenological description in terms of a three-quark core that dominates at high $Q^{2}$ and MB contributions that can make important contributions at lower $Q^{2}$. From the direct measurement of the MB contributions on the proton, and from measurements on the neutron at the photon point subtracted of the LF RQM predictions, one finds that for this resonance the MB contributions are of isovector nature, as shown in Fig. 5.16.